OpenStudy (anonymous):

Add three and one-eighth plus four and five-eighths. one and one-half seven and three-fourths seven and one-half one and three-fourths

OpenStudy (anonymous):

this is it @Hero @angel12310 @halorazer

OpenStudy (anonymous):

@angel12310

OpenStudy (anonymous):

oh and can we make this one quick? I really have to go

OpenStudy (anonymous):

3 1/8 + 4 5/8 (3+4) as the numbers on the side, (5+1)/8 =6/8=3/4 Now put them together . 7 3/4 So your answer is

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

okay\ thank you @halorazer !